Graph the line -4x+5y=20
I forgot how to do this, could you help :)

Answer by sajankumar(7)

the answer is actualy i dont know how to graph in internet but the so theanswer is like thise.
-4x+5y=20
solve for x
-4x+5(0)=20 ****** replace y to (0) zero
-4x=20
-4x/-4=20/-4 ****** divide both side with -4
x=-5
point it to your graph at left side side or the negative side of x

solve for y
-4(0)+5y=20 ****** replace x to (0) zero
5y=20
5y/5=20/5 ****** divide both side with 5
y=4
point it to your graph at upper side side or the positive side of y